OpenStudy (anonymous):

a light body and a heavy body have same linear momentum.which one has greater K.E.?

OpenStudy (ash2326):

Momentum=\(m\times v\) K.E=\(\frac{1}{2} m\times v^2\) light body's momentum= heavy body's momentum => velocity of light body is greater than velocity of heavy body so tell which one has the higher KE, you have the relation

OpenStudy (anonymous):

M1>M2 and let p1 = p2 i.e when the momentums of the bodies are same then K.E1/ K.E2 = M2/M1 i.e the lighter body will have more K.E than the heavier body.